Circular No. 7751

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fw IN UGC 2116 T. Puckett and M. Marcus, Mountain Town, GA, report the discovery

of an apparent supernova (mag 18.2) on an unfiltered CCD frame taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Nov. 11.15 UT. SN 2001fw is located at α = 2^{h}38^{m}10^{s}.45, δ = +30^{o}50′54′′.6 (equinox 2000.0), which is 8′′.8 west and 1′′.5 north of the center of UGC 2116. The new object was confirmed on CCD frames taken on Nov. 12.19. SN 2001fw is not present on images taken by Puckett on Feb. 5, Sept. 26, Oct. 3, and 26 (limiting mag ∼ 20.5), or on Palomar Sky Survey images taken on 1993 Nov. 20, 1991 Oct. 5 (limiting mag ∼ 21.0), and 1954 Dec. 25 (limiting mag ∼ 19.7).

SUPERNOVAE 2001ej, 2001fx, AND 2001fy M. Papenkova and W. D. Li, University of California at Berkeley, re-

port observations by LOTOSS (cf. IAUC 7514) of SN 2001ej and two new apparent supernovae in unfiltered Katzman Automatic Imaging Telescope (KAIT) images:

SN  2001 UT  α_{2000}  δ_{2000}  Mag.  Offset
2001fx Nov. 8.2 23^{h}39^{m}32.^{s}72 −22^{o}24′46′′.3 18.2 8′′.0 E, 1′′.3 N
2001fy Nov. 7.2 22 08 54.03 +40 20 50.3 18.7 2′′.9 W, 1′′.5 S

Additional magnitude estimates from CCD images taken by KAIT and by M. Schwartz with the Tenagra Observatory 0.5-m telescope (the latter denoted by T): SN 2001ej in UGC 3829, Mar. 24.3 UT, [19.0; Oct. 23.5 (T), Nov. 1.5, and 10.5, ∼ 17.5. SN 2001fx in IC 5345, Oct. 14.2, [19.0; 20.2, (hint); 25.2, (hint); Nov. 1.2, (hint); Nov. 10.2, ∼ 18.3; 11.2, ∼ 18.3 (T). SN 2001fy in UGC 11922, Nov. 1.2, [19.5; 12.2, ∼ 18.4 (T).

COMET C/2000 WM_{1} (LINEAR) Visual m_{1} and coma-diameter estimates: Oct. 8.82 UT, 11.1, 2′.8 (R.

Bouma, Groningen, The Netherlands, 0.31-m reflector); 23.04, 10.0, 2′.6 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 28.14, 10.0, 3′.0 (N. Biver, Poigny, France, 0.26-m reflector); Nov. 3.72, 9.4, 3′ (A. Baransky, Kiev, Ukraine, 0.36-m reflector); 9.74, 8.2, 7′.5 (Bouma, 15×80 binoculars); 12.12, 8.3, 6′ (P. M. Raymundo, Salvador, Brazil, 11×80 binoculars).

2001 November 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green